Problem 1.78:

A ship leaves the island of Guam and sails a distance 300 km at an angle 40.0 ∘ north of west.

Part A: In which direction must it now head so that its resultant displacement will be 110 km directly east of Guam? (Express your answer as an angle measured south of east). theta = ________ degrees South of East.

Part B: How far must it sail so that its resultant displacement will be 110 km directly east of Guam?

______ km
Answer #1

first displacement :

d1 = 300 km , 40 degree north of west

d1x = X- component = - 300 Cos40 = - 229.81km               towards west

d1y = Y- component = 300 Sin40 = 192.8 km               towards north

Net displacement ::

dnet = 110 km

dnetx = 110 km    towards east

dnety = 0 km

Let the second displacement be d2 and d2x and d2y are its components

then Along X-direction

dnetx = d1x + d2x

110 = - 229.81 + d2x

d2x = 339.81 km

then Along Y-direction

dnety= d1y + d2y

0 = 192.8 + d2y

d2y = - 192.8 km

d2 = sqrt (d2x2 + d2y2)

d2 = sqrt (339.812 + (-192.8)2)

d2 = 390.7 km

direction :

= tan-1 (d2y / d2x) = tan-1 (192.8 /339.81) = 29.6 degree south of east

B)

displacement = d2 = 390.7 km